# Scaling Behaviors of LLM Reinforcement Learning Post-Training: An Empirical Study in Mathematical Reasoning

### v1

@@ 3.3 Scaling up Model Size
We train models of varying sizes to convergence and compare their final test loss. As shown in Figure 4(a), larger models consistently achieve lower loss, improving monotonically with scale. The curve deviates from a strict power law: smaller models show weaker gains, suggesting diminishing returns at low parameter counts. A likely reason is that larger models inherit richer pre-trained representations, which reinforcement fine-tuning exploits for greater improvements than parameter growth alone would predict. Figure 4(b) further shows that as RL training progresses, larger models generate longer responses. This correlates with higher accuracy, indicating greater test-time scaling efficiency: additional inference tokens yield larger gains in bigger models.
We also benchmark our RL-tuned Qwen2.5 models (Qwen et al., 2025) against state-of-the-art open-source reasoning systems, including Qwen3 (Yang et al., 2025) and GPT-OSS (OpenAI et al., 2025). On our held-out set, the 7B and 14B models match or surpass dense Qwen3 counterparts of similar size, highlighting the effectiveness of RL post-training. Mixture-of-experts models such as Qwen3 and GPT-OSS achieve lower loss at much larger scales (20B–235B), with GPT-OSS-120B currently leading. These comparisons suggest that scaling beyond 14B will be necessary to fully characterize post-training behavior and compete with frontier MoE systems.
(a) Relation between test loss and model size NN when trained to convergence on sufficiently large data.
(b) Relation between response length and data size for instruct models.
Figure 4: Analysis of model scaling properties.
@@ 3.4 Scaling with Constrained Data and Reuse
Observation 4 Performance in data-constrained settings is primarily determined by the total number of parameter update steps ($S$). For a fixed $S$, the final test loss is remarkably insensitive to the data reuse factor ($\tau$), with no significant degradation up to $\tau=25$.
@@
We further consider the data-reuse scenario (Eq. 3), where high-quality data is limited but can be revisited multiple times. To simulate this, we partition the training set into smaller subsets while preserving the difficulty distribution (Details provided in Appendix A.4). Each subset is cycled through multiple times, with the reuse factor $\tau$ controlling how often each unique example is revisited. The total number of update steps $S$ is fixed across runs, and curriculum ordering is maintained so that problems are always presented from easy to hard. This ensures that performance differences arise solely from the degree of data reuse, rather than distributional or scheduling artifacts.
As shown in Figure 5, performance remains nearly unchanged for $\tau\leq 25$, while moderate degradation appears as $\tau$ increases further. At $\tau=100$, we observe clear signs of overfitting, indicating that repeated reuse eventually harms generalization. Overall, these results suggest that final performance is primarily governed by the total number of optimization steps rather than sample uniqueness, and that moderate data reuse is an effective strategy for RL fine-tuning with limited datasets.
@@ 3.5 Domain Transfer
(a) A consistent decrease in test loss is observed on various in-domain mathematics tasks.
(b) Mixed results are observed on out-of-domain tasks.
Figure 6: The effect of domain transfer, illustrated with the Qwen2.5-14B-Base model.
@@
In-Domain Generalization. Figure 6(a) shows consistent improvements on unseen mathematics tasks outside the training set. On benchmarks, from easy to hard, including GSM8K, MATH-500, AMC2023, AIME2024, test loss steadily decreases with training compute, suggesting that RL post-training enhances transferable reasoning skills within mathematics.
Out-of-Domain Generalization. As shown in Figure 6(b), results on OOD tasks are markedly different. For code generation (HumanEval) and STEM problems (SuperGPQA), performance gains marginally, indicating that RL fine-tuning is highly specialized. More strikingly, on logical reasoning (zebra_puzzle), the 14B model’s test loss increases during training, suggesting that intensive domain-specific optimization can negatively impact distinct reasoning skills.
3.6 Ablation on GRPO Hyperparameters
Observation 6 A larger GRPO rollout group size ($G$) is consistently more data-efficient, while the compute-optimal $G$ grows with the total computational budget.
Data-centric View. Figure 7(b) and 7(d) shows that larger rollout sizes consistently yield better sample efficiency: $G=32$ achieves the lowest test loss for the same number of unique samples. This supports the intuition that more responses per question provide a stronger advantage estimate and thus more effective gradient updates.
(a) 7B-Base: Loss vs. Compute
(b) 7B-Base: Loss vs. Data Size
(c) 7B-Instruct: Loss vs. Compute
(d) 7B-Instruct: Loss vs. Data Size
Figure 7: Effects of GRPO rollout size on training efficiency: (a–b) show compute/data scaling for 7B-Base; (c–d) for 7B-Instruct.
@@ 3.6 Ablation on GRPO Hyperparameters
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size $N$, data $D$, and compute $C$ (Kaplan et al., 2020), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (Hoffmann et al., 2022). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (Pearce & Song, 2024; Porian et al., 2024), while data-centric refinements examine pruning efficiency (Sorscher et al., 2022), repetition effects (Hernandez et al., 2022), gzip-based complexity predictors (Pandey, 2024), constrained or synthetic regimes (Muennighoff et al., 2023; Qin et al., 2025), and task transfer (e.g., translation) (Isik et al., 2024). Test-time compute amplification supplies an inference analogue to classical training laws (Snell et al., 2024).
RL post-training in LLMs. In RL, power-law trends similarly link capacity, interaction compute, and performance (Hilton et al., 2023); scaling RFT across horizon and compute improves mathematical and coding reasoning (DeepSeek-AI, 2025; Kimi Team, 2025; Mai et al., 2025b; Zhang et al., 2025a; b), while extended schedules (Liu et al., 2025), ultra-low-shot or single-example RL (Wang et al., 2025), and minimal-data efficiency paradigms (Li et al., 2025b) probe data–compute tradeoffs. Instability and uneven gains highlight fragile optimization (Zeng et al., 2025a; Yue et al., 2025), and multi-domain mixtures reveal both synergy and interference across math, code, and logic (Li et al., 2025c; Cheng et al., 2025).
Mathematical Reasoning with LLMs. Mathematical reasoning amplifies these dynamics: accuracy generally scales upward while verification behaviors remain inconsistent (Touvron et al., 2023); corpus volume and quality jointly shape attainable curves (Ye et al., 2024); multi-task math-generalist training diverges from specialist scaling trajectories (Yue et al., 2023); and RL with code execution induces additional behaviors such as emergent tool use concentrated in math problem solving (Zeng et al., 2025b). Collectively, evidence indicates that reasoning performance is governed by interacting axes of model size, data distribution/quality, training (supervised vs. RL) paradigm, and allocation of both training and inference compute, while unified laws for mathematical reasoning remain only partially characterized.
@@ 5 Discussion
Scaling Dependence on Evaluation Environment and Metrics. Reinforcement learning optimizes directly for environment rewards (Sutton & Barto, 2018), which in principle allows unbounded capability—as demonstrated by AlphaZero mastering board games (Silver et al., 2017), AlphaFold predicting protein structures (Jumper et al., 2021), and frontier LLMs such as Gemini-2.5-Pro achieving IMO-level performance (Huang & Yang, 2025). In contrast, text-based LLMs lack a well-defined RL environment, forcing us to rely on human-curated datasets as proxies. Test loss thus serves as a pragmatic but imperfect metric: it is monotonic and convergent, yet heavily dependent on dataset construction and task difficulty, with different benchmarks (e.g., GSM8K vs. AIME, Section 3.5) showing distinct convergence rates. This task dependence makes the absolute coefficients of our fitted scaling laws ($k(N),E$) difficult to interpret universally. Prior work proposed “intrinsic performance”—the minimum compute needed to reach a target reward—as a normalization across environments (Hilton et al., 2023), but we did not find an analogous measure in large-scale LLMs. Establishing principled, environment-independent evaluation protocols remains an open and critical challenge for RL-based scaling studies.
Scaling Dependence on Model Scale. Our study of models from 0.5B to 14B parameters shows that larger models exhibit greater sample and compute efficiency in RL post-training. This parameter range, however, is still below that of frontier models. We anticipate these advantages may not extend indefinitely, as performance can saturate on simple benchmarks like GSM8K and fail to converge on extremely hard problems. A plausible hypothesis is the existence of a ”sweet spot” or critical model scale—large enough for efficient RL adaptation but not so over-parameterized that returns diminish. Identifying this optimal scale is still an important direction for future work.
@@
Dependence on RL Algorithm. Our analysis is based on GRPO, a mainstream and stable RL post-training algorithm for LLMs that uses an actor-only design and normalizes rewards across responses. Comparative study with alternative RL algorithms  (Cui et al., 2025) reports minor differences in training curves. Whether more advanced algorithms can significantly improve sample efficiency or stability—and thereby reshape the scaling frontier—remains an important open question.
@@
Future of LLM Agent. The integration of reinforcement learning with agentic LLMs is increasingly viewed as a promising direction (Zhang et al., 2025a; b). Both theoretical and empirical studies show that augmentations such as external tool use and long-term memory can substantially boost model performance (Lin & Xu, 2025; Houliston et al., 2025; Mai et al., 2025a). We anticipate that such agentic mechanisms will markedly improve the scaling behavior of RL-trained LLMs: by offloading deterministic computations to tools and focusing learning on high-level decision making, these models could achieve much higher efficiency, effectively shifting the performance frontier upward for a given compute or data budget. Understanding the scaling laws of these agentic systems is, therefore, a key and exciting avenue for future research.
@@ 6 Conclusion
7 Statement
7.1 Ethics statement
The authors have read and adhered to the ICLR Code of Ethics. This work is foundational in nature, focusing on the scaling properties of large language models in the domain of mathematical reasoning. Our research exclusively utilizes publicly available and previously published resources, including open-source models (e.g., Qwen2.5) and established datasets (e.g.,guru-RL-92k), thereby mitigating concerns related to data privacy, human subjects, or the release of sensitive information. The application domain of mathematical problem-solving does not inherently present risks of direct societal harm. The primary ethical consideration associated with this work is the environmental impact of the computational resources required for large-scale model training, a challenge common to the field. We believe that by providing insights into efficient resource allocation, our work contributes positively to mitigating this concern for future research.
7.2 Reproducibility statement
We are committed to ensuring the reproducibility of our research. All aspects of our experimental design are detailed to facilitate this. The specific models used (Qwen2.5 series), are described in Section 2. The training and evaluation datasets, including our held-out set, are also detailed in Section 2, with a full breakdown provided in Appendix A.1. Our reinforcement learning methodology is based on the GRPO algorithm, and all relevant hyperparameters used for training and evaluation are listed in Appendix A.2. The evaluation protocol, including the precise definition of our Test Loss metric and the prompt templates for all tasks, is outlined in Section 2 and Appendix A.3. All figures and observations presented are directly generated from the raw data collected during our experiments. At last, we provide the experiment source code in Supplementary Material along with the submission.
@@ A.4 Data Reuse Experiment Setup
To systematically evaluate the effect of data reuse under constrained data scenarios, we design controlled experiments where all runs are trained with the same total data size but different levels of data repetition. Each run randomly samples a subset from the full training corpus and repeats this subset sufficiently many times to exactly match the target data budget (i.e., $\text{subset size}\times\tau=\text{total data size}$). Unlike Muennighoff et al. (2023), subsets are sampled independently for each run rather than sampling within the larger subsets, to mitigate sampling bias and balance stochasticity across conditions. To remain consistent with the Curriculum Learning setting of the main experiments, examples within each subset are ordered by increasing difficulty; across epochs, this difficulty schedule is preserved and repeated rather than reshuffled, as illustrated in Figure 8.
@@ B.1 Performance on In-Domain and Out-of-Domain Tasks
Figure 10: Test loss on in-domain and out-of-domain benchmarks vs data size for Instruct models. It shows modest positive transfer on in-domain tasks, with limited or negative transfer on OOD tasks.
B.2 Peformance Compared with Advance Models
@@ C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters ($N$) and the number of processed tokens ($T$). The costs for the fundamental operations are:
@@ C.2 Coefficient Comparison
Figure 11: Fitted coefficients $k(N)$ and $E_{D}(N)$ comparison over $L(N,D)$ and $L(N,C)$. Upper: base model. Bottom: instruct model.
@@
The observation from Figure 11 also matches with this conclusion.
@@ C.3 Fitting for K and E
Table 7: $L(N,C)$ Fitting Results: $\log\text{L}=-k_{C}\log C+E_{C}$
Model Size Base Instruct
$k_{C}$ $E_{C}$ $R_{C}^{2}$ $k_{C}$ $E_{C}$ $R_{C}^{2}$
0.5B 0.0075 0.1140 0.485 0.0055 0.0744 0.313
1.5B 0.0338 0.5889 0.813 0.0063 0.0718 0.293
3B 0.0168 0.2448 0.621 0.0153 0.2013 0.570
7B 0.0370 0.5562 0.789 0.0436 0.6600 0.809
14B 0.0680 1.1023 0.921 0.0454 0.6361 0.817
Table 8: $L(N,D)$ Fitting Results: $\log\text{TestLoss}=-k_{D}\log D+E_{D}$
Model Size Base Instruct
$k_{D}$ $E_{D}$ $R_{D}^{2}$ $k_{D}$ $E_{D}$ $R_{D}^{2}$
0.5B 0.0070 0.0117 0.483 0.0054 0.0002 0.314
1.5B 0.0268 0.0839 0.796 0.0065 -0.0150 0.294
3B 0.0181 0.0127 0.617 0.0164 -0.0112 0.573
7B 0.0409 0.0359 0.792 0.0469 0.0399 0.812
14B 0.0739 0.1200 0.921 0.0504 -0.0167 0.818
Appendix D Response Length

### v2

@@ 3.3 Scaling up Model Size
(a) Fig (a) shows Relation between test loss and model size NN for our trained model and SOTA models demonstrates the effectiveness of our training.
(b) Fig (b) shows relation between response length and data size for instruct models,larger models generate longer responses except for 32B model.
Figure 8: Analysis of model scaling properties shows the effectiveness of our training process in this empirical study.
We train models of varying sizes to convergence and compare their final test loss. As shown in Figure 8(a), larger models consistently achieve lower loss, improving monotonically with scale. The curve deviates from a strict power law: smaller models show weaker gains, suggesting diminishing returns at low parameter counts. A likely reason is that larger models inherit richer pre-trained representations, which reinforcement fine-tuning exploits for greater improvements than parameter growth alone would predict. Figure 8(b) further shows that as RL training progresses, larger models generate longer responses except for 32B model. This correlates with higher accuracy, indicating greater test-time scaling efficiency: additional inference tokens yield larger gains in bigger models.
We also benchmark our RL-tuned Qwen2.5 models (qwen2025qwen25technicalreport) against state-of-the-art open-source reasoning systems, including Qwen3 (yang2025qwen3technicalreport) and GPT-OSS (openai2025gptoss120bgptoss20bmodel), detailed in Table 4. On our held-out set, the 32B and 72B models match or surpass dense Qwen3 counterparts of similar size, highlighting the effectiveness of RL post-training. Mixture-of-experts models such as Qwen3 and GPT-OSS achieve approximate loss at much larger scales (235B), with GPT-OSS-120B currently leading. These comparisons suggest that scaling across 0.5B-72B will be necessary to fully characterize post-training behavior and compete with frontier MoE systems.
@@ 3.4 Scaling with Constrained Data and Reuse
When unique data is scarce, a critical question is whether repeating data is effective. We investigate this Data Reuse Scenario by fixing the total data budget and varying the reuse factor $\tau$. Specifically, we aim to identify the optimal reuse factor $\tau$ that minimizes the test loss:
$\underset{\tau}{\text{argmin}};L(\tau)\quad\text{s.t.}\quad D_{\text{unique}}\times\tau=D_{\text{total}},$ (9)
Observation 4 Performance in data-constrained settings is primarily determined by the total number of used data ($D_{\text{total}}$). For a fixed $D_{\text{total}}$, the final test loss is remarkably insensitive to the data reuse factor ($\tau$), with no significant degradation up to $\tau=25$.
@@
We further consider the data-reuse scenario, where high-quality data is limited but can be revisited multiple times. To simulate this, we partition the training set into smaller subsets while preserving the difficulty distribution (Details provided in Appendix A.4). Each subset is cycled through multiple times, with the reuse factor $\tau$ controlling how often each unique example is revisited. The total number of used data $D_{\text{total}}$ is fixed across runs, and curriculum ordering is maintained so that problems are always presented from easy to hard. This ensures that performance differences arise solely from the degree of data reuse, rather than distributional or scheduling artifacts.
As shown in Figure 9, performance remains nearly unchanged for $\tau\leq 25$, while moderate degradation appears as $\tau$ increases further. At $\tau=100$, we observe clear signs of overfitting, indicating that repeated reuse eventually harms generalization. Overall, these results suggest that final performance is primarily governed by the total number of used data rather than sample uniqueness, and that moderate data reuse is an effective strategy for RL fine-tuning with limited datasets.
@@ 3.5 Domain Transfer
Figure 10: The effect of domain transfer, illustrated with the Qwen2.5-72B-Instruct model.
@@
In-Domain Generalization. Figure 10 shows consistent improvements on unseen mathematics tasks outside the training set. On benchmarks, from easy to hard, including GSM8K, MATH-500, AMC2023, AIME2024, test loss steadily decreases with training compute, suggesting that RL post-training enhances transferable reasoning skills within mathematics.
Out-of-Domain Generalization. As shown in Figure 10, results on OOD tasks are markedly different. For code generation (HumanEval) and STEM problems (SuperGPQA), performance gains marginally, indicating that RL fine-tuning is highly specialized.On logical reasoning (zebra_puzzle), performance degrades for larger models, suggesting that intensive optimization on mathematical reasoning may interfere with or "damage" other distinct reasoning abilities.
@@
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size NN, data $D$, and compute $C$ (kaplan2020scaling), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (hoffmann2022training). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (pearce2024reconciling; porian2024resolving), while data-centric refinements examine pruning efficiency (sorscher2022beyond), repetition effects (hernandez2022scaling), gzip-based complexity predictors (pandey2024gzip), constrained or synthetic regimes (muennighoff2023scaling; qin2025scaling), and task transfer (e.g., translation) (isik2024scaling). Test-time compute amplification supplies an inference analogue to classical training laws (snell2024scaling).
RL post-training in LLMs. In RL, power-law trends similarly link capacity, interaction compute, and performance (hilton2023scaling); scaling RFT across horizon and compute improves mathematical and coding reasoning (guo2025deepseekr1; kimi2025scaling; mai2025agent; zhang2025landscapeagenticreinforcementlearning; zhang2025surveyreinforcementlearninglarge), while extended schedules (liu2025prorl), ultra-low-shot or single-example RL (wang2025reinforcement), and minimal-data efficiency paradigms (li2025limr) probe data–compute tradeoffs. Instability and uneven gains highlight fragile optimization (zeng2025simplerlzoo; yue2025doesrlincentivize), and multi-domain mixtures reveal both synergy and interference across math, code, and logic (li2025onedomainhelps; cheng2025revisiting).
Mathematical Reasoning with LLMs. Mathematical reasoning amplifies these dynamics: accuracy generally scales upward while verification behaviors remain inconsistent (touvron2023scaling); corpus volume and quality jointly shape attainable curves (ye2024skywork); multi-task math-generalist training diverges from specialist scaling trajectories (yue2023mammoth); and RL with code execution induces additional behaviors such as emergent tool use concentrated in math problem solving (zeng2025agent). Collectively, evidence indicates that reasoning performance is governed by interacting axes of model size, data distribution/quality, training (supervised vs. RL) paradigm, and allocation of both training and inference compute, while unified laws for mathematical reasoning remain only partially characterized.
@@ 5 Discussion
Scaling Dependence on Evaluation Environment and Metrics. Reinforcement learning optimizes directly for environment rewards (10.5555/3312046), which in principle allows unbounded capability—as demonstrated by AlphaZero mastering board games (silver2017masteringchessshogiselfplay), AlphaFold predicting protein structures (jumper2021highly), and frontier LLMs such as Gemini-2.5-Pro achieving IMO-level performance (huang2025gemini25procapable). In contrast, text-based LLMs lack a well-defined RL environment, forcing us to rely on human-curated datasets as proxies. Test loss thus serves as a pragmatic but imperfect metric: it is monotonic and convergent, yet heavily dependent on dataset construction and task difficulty, with different benchmarks (e.g., GSM8K vs. AIME, Section 3.5) showing distinct convergence rates. This task dependence makes the absolute coefficients of our fitted scaling laws ($k(N),E$) difficult to interpret universally. Prior work proposed “intrinsic performance”—the minimum compute needed to reach a target reward—as a normalization across environments (hilton2023scaling), but we did not find an analogous measure in large-scale LLMs. Establishing principled, environment-independent evaluation protocols remains an open and critical challenge for RL-based scaling studies.
Scaling Dependence on Model Scale. Our study of models from 0.5B to 72B parameters shows that larger models exhibit greater sample and compute efficiency in RL post-training. This parameter range allows us to characterize the scaling limits. We found that these advantages do not extend indefinitely. Our analytic learning efficiency term $k(N)$ in Eq.6 and Eq.8, explicitly confirms that the efficiency gains follow a saturation curve toward a limit ($K_{\max}$). This finding implies that scaling up models beyond a certain point, while still yielding absolute performance gains, suffers from diminishing marginal returns in efficiency.
@@
Dependence on RL Algorithm. Our analysis is based on GRPO, a mainstream and stable RL post-training algorithm for LLMs that uses an actor-only design and normalizes rewards across responses. Comparative study with alternative RL algorithms  (cui2025entropymechanismreinforcementlearning) reports minor differences in training curves. Whether more advanced algorithms can significantly improve sample efficiency or stability—and thereby reshape the scaling frontier—remains an important open question.
@@
Future of LLM Agent. The integration of reinforcement learning with agentic LLMs is increasingly viewed as a promising direction (zhang2025landscapeagenticreinforcementlearning; zhang2025surveyreinforcementlearninglarge). Both theoretical and empirical studies show that augmentations such as external tool use and long-term memory can substantially boost model performance (lin2025understandingtoolintegratedreasoning; houliston2025provablebenefitsintoollearning; mai2025agentrlscalinglaw). We anticipate that such agentic mechanisms will markedly improve the scaling behavior of RL-trained LLMs: by offloading deterministic computations to tools and focusing learning on high-level decision making, these models could achieve much higher efficiency, effectively shifting the performance frontier upward for a given compute or data budget. Understanding the scaling laws of these agentic systems is, therefore, a key and exciting avenue for future research.
@@ A.4 Data Reuse Experiment Setup
To systematically evaluate the effect of data reuse under constrained data scenarios, we design controlled experiments where all runs are trained with the same total data size but different levels of data repetition. Each run randomly samples a subset from the full training corpus and repeats this subset sufficiently many times to exactly match the target data budget (i.e., $\text{subset size}\times\tau=\text{total data size}$). Unlike muennighoff2023scaling, subsets are sampled independently for each run rather than sampling within the larger subsets, to mitigate sampling bias and balance stochasticity across conditions. To remain consistent with the Curriculum Learning setting of the main experiments, examples within each subset are ordered by increasing difficulty; across epochs, this difficulty schedule is preserved and repeated rather than reshuffled, as illustrated in Figure 11.
@@ B.1 Performance on In-Domain and Out-of-Domain Tasks
Figure 13: Test loss on in-domain and out-of-domain benchmarks vs data size for Instruct models.
B.2 Ablation on GRPO Hyperparameters
(a) 7B-Base: Loss vs. Compute
(b) 7B-Base: Loss vs. Data Size
(c) 7B-Instruct: Loss vs. Compute
(d) 7B-Instruct: Loss vs. Data Size
Figure 14: Effects of GRPO rollout size on training efficiency
Data-centric View. Figure 14(b) and 14(d) shows that larger rollout sizes consistently yield better sample efficiency: $G=32$ achieves the lowest test loss for the same number of unique samples. This supports the intuition that more responses per question provide a stronger advantage estimate and thus more effective gradient updates.
B.3 Peformance Compared with Advance Models
@@ C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters (NN) and the number of processed tokens ($T$). The costs for the fundamental operations are:
@@ C.2 Coefficient Comparison
The observation from Figure 4 and Figure 7 also matches with this conclusion.
@@ C.3 Fitting for K and E
Table 5: Uncertainty Analysis for raw data: Base and Instruct Models (Holdout Score)
Model Base Instruct
Test Loss Avg Std SEM Test Loss Avg Std SEM
0.5B 0.9419 0.0082 0.0048 0.9458 0.0073 0.0042
1B 0.9129 0.0091 0.0053 0.8988 0.0098 0.0057
3B 0.8582 0.0129 0.0074 0.8281 0.0112 0.0065
7B 0.7148 0.0147 0.0085 0.6777 0.0142 0.0082
14B 0.6051 0.0149 0.0086 0.5588 0.0143 0.0083
32B 0.4937 0.0056 0.0032 0.4579 0.0127 0.0073
72B 0.4359 0.0143 0.0082 0.4320 0.0140 0.0081
Table 6: Comparison of $k_{max}$ and $N_{0}$ Parameters Across Fitting Scenarios
Source Metric Scenario $k_{max}$ $N_{0}$ (B) $R^{2}$
Base L(N,C) Intra-model 0.1349 13.09 0.9955
L(N,C) Inter-model 0.1518 17.37 0.9944
L(N,D) Intra-model 0.1348 11.52 0.9953
L(N,D) Inter-model 0.1631 16.95 0.9947
Instruct L(N,C) Intra-model 0.1276 17.27 0.9970
L(N,C) Inter-model 0.1443 28.33 0.9950
L(N,D) Intra-model 0.1325 17.08 0.9970
L(N,D) Inter-model 0.1484 27.15 0.9949
Appendix D A Loss Decomposition Model for Scaling Analysis
During the analysis, we found a more generalized form of the potential scaling law function that fits the curves well. This model fits the same dataset as the main experiments and is included here as a formally documented alternative for future research.
D.1 Loss Decomposition Model
The General Loss Decomposition.
We construct the generalized formula as follows, based on the observation of the loss composition of post-training and the experiment data:
$L(N,D)=L_{\infty}+G(N)+\lambda(N)\cdot P(N,D)$ (16)
Each term in Equation equation 16 represents a clear part decomposing the loss:
  • $L_{\infty}$ denotes the irreducible loss, representing the fundamental loss floor that persists even with infinite model capacity and unlimited data. It reflects task-intrinsic uncertainty and noise that cannot be eliminated by improved modeling or additional training, such as inherent stochasticity in the environment or irreducible mismatch between training and evaluation distributions.
  • $G(N)$ denotes the model-limited loss, capturing the asymptotic loss floor imposed by finite model capacity NN in the limit of infinite data. It corresponds to the capacity-dependent performance frontier of models with size NN.
  • $\lambda(N)$ denotes the learnable capacity, defined as the maximum achievable reduction in loss that a model of size NN can attain through post-training, beyond its model-limited loss. In RL post-training settings, this term can depend on the pretraining regime, as well as the degree of mismatch between the pretraining and post-training task settings. Conceptually, the learnable capacity is governed by two opposing effects: larger models generally have greater capacity to extract and acquire new knowledge, while simultaneously having already absorbed more information during pretraining, leaving less headroom for additional improvement via RL. As a result, the monotonic dependence of $\lambda(N)$ on model size is non-trivial, and its precise modeling likely requires additional assumptions and empirical characterization.
  • $P(N,D)$ denotes the learning progress, a normalized function taking values in $[0,1]$ that quantifies the fraction of the learnable capacity $\lambda(N)$ realized when training on a dataset of size $D$.
Instantiated Model.
Following prior empirical scaling law studies, we parameterize the model-limited loss as
$G(N)=\left(\frac{N_{0}}{N}\right)^{\alpha},$ (17)
reflecting the observed power-law dependence of loss on model size in the infinite-data regime (kaplan2020scaling).
Empirically, learning curves exhibit an S-shaped transition when plotted in log–log coordinates (e.g. Figure 1). Motivated by this observation, we model the learning progress term $P(N,D)$ as a logistic function in $\log D$,
$P(N,D)=\frac{1}{1+\left(\frac{D}{D_{0}(N)}\right)^{\beta}},$ (18)
where $D_{0}(N)$ denotes the characteristic dataset scale at which half of the learnable capacity is realized, and is treated as a NN-dependent parameter.
Combining the above components, we arrive at the following instantiated loss model:
$L(N,D)=L_{\infty}+\left(\frac{N_{0}}{N}\right)^{\alpha}+\frac{\lambda(N)}{1+\left(\frac{D}{D_{0}(N)}\right)^{\beta}}$ (19)
This is what we used to fit and extrapolation in Figure 15
D.2 Predictability and Extrapolation
(a) Intra-model prediction using the early 30% of training steps.
(b) Inter-model extrapolation for 72B under fixed shared shape.
Figure 15: Extrapolation of the loss decomposition model (Equation equation 19). (a) Intra-model prediction using partial learning-curve observations. (b) Inter-model extrapolation for 72B with global exponents fixed from smaller models.
We evaluate the extrapolation behavior of the loss–decomposition model (Equation equation 19) by applying it to the same experimental learning-curve data used in the main analysis, and by testing its performance under two complementary settings (Figure 15).
Intra-model prediction.
Using only the first 30% of training steps for each model, the fitted curves closely match the held-out portions(Figure 15a). This indicates that the internal S-shaped structure of the model provides a sufficiently strong inductive bias for completing a single learning curve from early observations. See Table 7 for detailed fitting result.
Inter-model extrapolation.
We fit all global exponents and the model-limited term using models up to 32B, and then extrapolate the resulting shared functional form to 72B by calibrating only two model-specific parameters, $\lambda(72\text{B})$ and $D_{0}(72\text{B})$. The extrapolated curve aligns well with the observed 72B trajectory across the full data range, reflecting that the functional shape inferred from smaller models remains compatible with larger-scale behavior under this light calibration.
The fit is reasonably strong, indicating that the proposed formulation may capture key structural tendencies of the underlying scaling behavior.
Table 7: Fitting details for $L(N,D)$ using only 30% of the training steps for each model.
Model Size Base
$D_{0}$ lambdalambda $R^{2}$
3B 5109.6316 0.0734 0.995
7B 3725.6374 0.1882 0.995
14B 4554.7619 0.2520 0.995
32B 3576.0323 0.3279 0.995
72B 5861.6228 0.3084 0.995
D.3 Discussion: Effective log–log slope
To relate the loss decomposition model to the slope-based formulation used in the main text, we examine the local behavior of $L(N,D)$ in log–log coordinates with respect to the data scale $D$. Specifically, we define the effective slope
$k(N,D)\;:=\;-\,\frac{\partial\log L(N,D)}{\partial\log D},$
which corresponds to the exponent in a local power-law approximation of the form $\log L\approx-k\log D+\text{const}$.
For the loss decomposition model (Equation equation 19), the induced $k(N,D)$ is a smooth function of $D$ that vanishes in both the low-data and high-data limits, and attains its maximum around the characteristic scale $D\approx D_{0}(N)$. Evaluating the slope at this point yields a natural definition of the maximal effective slope,
$k_{\max}(N)=\frac{K_{\max}}{1+S(N)},\qquad K_{\max}:=\frac{\beta}{2},\quad S(N):=\frac{2\!\left(L_{\infty}+(N_{0}/N)^{\alpha}\right)}{\lambda(N)}.$ (20)
The resulting $k_{\max}(N)$ depends only on the model size NN through the parameters of the loss decomposition, and is uniformly bounded by $K_{\max}=\beta/2$.
From this perspective, the slope function $k(N)$ adopted in the main analysis (Equation equation 2) can be interpreted as a parsimonious, low-parameter approximation to the effective maximal slope $k_{\max}(N)$ in Equation equation 20. This establishes structural consistency between the two descriptions: the loss decomposition model potentially captures the finer-grained $(N,D)$-dependent behavior, while the main-text formulation summarizes its dominant NN-dependent trend in a compact form.
D.4 Conclusion
The loss decomposition model captures key empirical characteristics of RL post-training scaling behavior.
However, several components remain underdetermined, including the construction of the learnable capacity $\lambda(N)$, the dependence of the characteristic dataset scale $D_{0}(N)$ on model size, the role of the pretraining process, and the impact of mismatches between pretraining and post-training task settings.
For these reasons, we present this model as an appendix-level discussion rather than a core component of the main text. We encourage future work to build on this formulation to further investigate and refine scaling laws for LLM-based reinforcement learning.
Appendix E Response Length